OpenStudy (horsegirl27):

MEDAL AND FAN Please help me out!!! I'm really confused and I have to get this done quickly!!!!!

OpenStudy (horsegirl27):

A scientist is studying the growth of a particular species of plant. He writes the following equation to show the height of the plant f(n), in cm, after n days: f(n) = 8(1.05)n Part A: When the scientist concluded his study, the height of the plant was approximately 11.26 cm. What is a reasonable domain to plot the growth function? (4 points) Part B: What does the y-intercept of the graph of the function f(n) represent? (2 points) Part C: What is the average rate of change of the function f(n) from n = 2 to n = 6, and what does it represent? (4 points)

OpenStudy (horsegirl27):

Part B: The y-intercept represents the height of the plant at the beginning. That's all I have so far. Is that right?

OpenStudy (horsegirl27):

part A: ok, this is stupid, i just can't think straight, soooo plz explain a domain to me. part C: do i just have to solve it for both 2 and 6?

OpenStudy (haseeb96):

the equation is f(n)=8(1.50)n now f(n) is the height of plant while n is days in part A the height f(n) is given that is f(n)=11.26cm now put it in the above equation at the place of f(n) 11.26 = 8(1.50)n now solve it for n

OpenStudy (haseeb96):

and show me what did u get the value of n?

OpenStudy (horsegirl27):

ok

OpenStudy (horsegirl27):

I multiply 8 by 1.5 and I get 11.26=12n, then divide by 11.26 to get 1.07=n or i could round even more and get 1.1=n

OpenStudy (haseeb96):

no n= 0.938 here n is the domain so the reasonable domain is 0.938 day

OpenStudy (horsegirl27):

ok

OpenStudy (horsegirl27):

how did u get .938?

OpenStudy (horsegirl27):

Thanks for explaining part A. Now, am i right about what I have to do for C? can u check my work when i do it to make sure i did it right?

OpenStudy (haseeb96):

11.26/8(1.50)= n now solve it

OpenStudy (horsegirl27):

wouldn't it just be 11.26/12=n?

OpenStudy (haseeb96):

now divide them by the help of calculator

OpenStudy (horsegirl27):

ok .938

OpenStudy (mathmath333):

\(11.26=8\times 1.05 \times n\\n=\dfrac{11.26}{8.4}\)

OpenStudy (mathmath333):

\(\cancel{=\dfrac{11.26}{12}}\)

OpenStudy (haseeb96):

average rate = f(b)-f(a)/b-a here a =2 and b= 6 now put f(2)= 8(1.50)(2) = 24 f(6) = 8(1.50)(6)=72 according to the average rate formula 72-24/6-2 12 is the average rate of function from n=2 to n=6

OpenStudy (horsegirl27):

ok, thank you :)

OpenStudy (mathmath333):

in the question it is \(f(n) = 8(\color{green}{1.05})n\) not \(f(n) = 8(\color{red}{1.50})n\)

OpenStudy (horsegirl27):

oh ok

OpenStudy (haseeb96):

sorry i did a small mistake can u correct it now

OpenStudy (horsegirl27):

yeah i think

OpenStudy (anonymous):

horsegirl, your function isn't right. "n" is an exponent on that test, I remember back when i took it.

OpenStudy (horsegirl27):

oh ok

OpenStudy (horsegirl27):

can u help please?

OpenStudy (anonymous):

For A, you need to figure out how many days it will take to reach 11.26cm, which just means make a table and plug in numbers until it works. I'm only up in mod 7, which teacher do you have?

OpenStudy (horsegirl27):

I'm actually in 6, but retaking this

OpenStudy (horsegirl27):

well i go to lss, i have ms. estes